Find the mean of a random variable z? z 0 1 2 3 f(z) 1/6 2/6 2/6 1/6


1
Expert's answer
2021-09-08T11:48:13-0400

E(Z) ="\sum z. f(z)"

="(0*\frac 16+1*\frac 2 6+2*\frac 2 6+3*\frac 1 6)"

="\frac 96"

=1.5